4(7x + 10) = 3(9x – 5) + 20
Answer by checkley71(8403)

4(7X+10)=3(9X-5)+20
28X+40=27X-15+20
28X-27X=-15+20-40
X=-35
PROOF
4(7*-35+10)=3(9*-35-5)+20
4(-245+10)=3(-315-5)+20
4*-235=3*-320+20
-940=-960+20
940=940